OpenStudy (anonymous):

The first four terms of a sequence are shown below. 7, 4, 1, -2 Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1

OpenStudy (anonymous):

Well 7-3=4, 4-3=1, 1-3=-2, therefore Its going to be f(n) -3.

OpenStudy (anonymous):

Thank you i was going to answer that but i wasnt sure :D